OpenStudy (anonymous):

Given the following two equations, what is the value of a? a + 2b =12 2a + b = 6 3 4 -3 6 0

OpenStudy (anonymous):

solve a for one of the equations and substitute that a into the other equation. Now you can find the value of b, and using that value for b in the first equation will give you the value of a. a + 2b = 12 a = 12 - 2b Substituting into second equation with a = 12 - 2b 2a + b = 6 thusly becomes -> 2(12-2b) = 6 24 - 4b = 6 -4b = 6 - 24 b = -18/-4 => b = 4.5 Now use this b in the first equation a + 2b = 12 a + 2*4.5 = 12 a = 12 - 9 => a = 3. You can use the b you found in anywhere you would like to find a. Hope this helped.

OpenStudy (zpupster):

a + 2b =12 2a + b = 6 give you 2 ways to solve first: get rid of b we do that by mult. 2nd equation by -2 so a + 2b =12 -2(2a + b = 6 ) we get a + 2b =12 -4a +-2b = -12 add the 2 equations 3a+0=0 so a = 0 and b would = 6 plugging them back makes the equation true second way: a + 2b =12 solve for a, subtract 2b both sides a=12-2b take this and sustitute in 2nd equation for a 2(12-2b)+ b = 6 24-4b+b=6 24-3b=6 subtract 24 both sides -3b=-18 b=6 plug this back in original equations

OpenStudy (anonymous):

Wow.. thanks guys